OpenStudy (impracticaljoker):

(04.03 LC) Which of the following options is an equivalent function to f(x) = 3(2)3x? (2 points) f(x) = 3(8)x f(x) = 24x f(x) = 27(8)x f(x) = 3(8x)

OpenStudy (princeharryyy):

? it's not clear. I mean the expression.

OpenStudy (mathmale):

Word to the wise: There are much better ways to indicate exponentiation, and you should definitely be using them. For example, 3(8)x literally means 3 times 8 times x and is not an exponential function at all. f(x)=3(8)x should be written as f(x)=3(8)^x, or \[f(x)=3(8)^x\]

OpenStudy (mathmale):

does this matter? You bet it does. Note how another student has already told you that your post is not clear to him / her.

OpenStudy (mathmale):

f(x) = 3(2)3x should be written as\[f(x)=3(2)^{3x}\]

OpenStudy (mathmale):

Try again. See if you can find the correct answer and show how you got it.